OpenStudy (anonymous):

If f(x) = x + 4, find f(3). If f(x) = −4x − 1, find f(−2). If g(x) = x2 + 3, find g(4). If h(x) = 2x + 1, find the value of x if h(x) = 11. If g(x) = 2(x − 4), find the value of x if g(x) = 20.

OpenStudy (anonymous):

For all of these, put the number in parentheses where x is and solve. First one: f(3)=3+4 f(3)=7

OpenStudy (anonymous):

Omg Your A Life Saver Thank You !

OpenStudy (anonymous):

The 4th and 5th one replace h(x) or g(x) with that number and solve for x. 4th one: 11=2x+1 10=2x 5=x